To what temperature must a given mass of nitrogen at 0⁰c be heated so that both its volume and pressure will be doubled?
1
Expert's answer
2020-06-29T14:14:36-0400

Let the initial volume of the gas "V_i=V"

Final volume of the gas"V_f=2V"

initial pressure of the gas"P_i=P"

final pressure of the gas "P_f=2P"

Initial temperature of the gas "T_i=0+273=273k"

Final temperature of the gas "T_f=?"

total number of mole of the gas is not getting change hence,

Using the ideal gas equation, total number of moles of the gas "n=\frac{PV}{RT}"

Hence,

"\frac{P_iV_i}{T_i}=\frac{P_fV_f}{T_f}"

Now, equating both side,

"\Rightarrow T_f=\frac{P_fV_fT_i}{P_iV_i}"

Now, substituting the values,

"\Rightarrow T_f=\frac{2P\times 2V\times 273k}{PV}"

"\Rightarrow T_f=4\times 273k=1092k"

"\Rightarrow T_f=1029-273=756^\circ C"